# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

### v5

@@ Abstract
Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge. In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies. By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF [1] and cosine similarity [2] to identify related episodes and enhance the overall story structure. Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
Index Terms:
Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human Feedback
**footnotetext: Equal contribution.footnotetext: Corresponding author.
I Introduction
Deep learning has transformed multiple domains including NLP, time series analysis and computer vision [3, 4, 5, 6, 7, 8]. Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [9]. However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge [10]. For instance, [11] pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
@@ I Introduction
  • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval.
  • We outperform baseline GPT model [16] in detecting continuity errors and maintaining overall narrative coherence.
II Method
@@
Figure 1: The SCORE framework for improving AI-generated story coherence. (a) Extracts key item statuses per episode. (b) Conducts detailed analysis and summaries of each episode. (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
II-A Continuity Analysis and Key Item Status Correction
@@
II-B Key Item Interaction Analysis
@@
II-C Similarity-Based Episode Evaluation and Sentiment Analysis
@@
For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores between the current episode or query and all other episodes in the vector space. Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$. The top $N$ episodes with the highest $S(e_{c},e_{p})$ scores are retrieved for further analysis, providing a relevant summary of episodes for evaluation or answering questions.
@@
TABLE I: Performance comparison of models with and without SCORE
Model Consistency Coherence Item Status Complex Question
GPT-4o-mini 78.2 76.7 0 24.56
+ SCORE 82.6 (↑4.4) 77.5 (↑0.8) 80.5 (↑80.5) 63.0 (↑38.44)
GPT-4o 86.78 82.21 0 76.32
+ SCORE 88.68 (↑1.9) 89.91 (↑7.7) 96 (↑96) 88.75 (↑12)
GPT-4 83.21 84.32 0 82.34
+ SCORE 85.61 (↑2.4) 86.9 (↑2.58) 98 (↑98) 89.45 (↑7.11)
III Experiments
@@
III-A Dataset Preparation
We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy. Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements. This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
III-B Baselines
@@
III-C Evaluation Process
For each episode in the dataset, the evaluation process involved several key steps. First, we conducted an initial evaluation by directly uploading the files to ChatGPT, testing if GPT could assess the story correctly without any prompts. We used GPT-4o-mini, GPT-4o, and GPT-4 for narrative evaluation.
Next, we configured the LLM for a more detailed evaluation, using our preprocessed files to ensure accurate tracking of key items across episodes. We also employed the Retrieval-Augmented Generation (RAG) framework to retrieve episodes similar to the one being evaluated. By using FAISS to calculate cosine similarity scores between episodes, the system was able to retrieve episodes that were semantically and emotionally aligned with the current one, providing additional context. This enriched context was then used to construct GPT prompts, enabling a more thorough evaluation of the episode.
III-D Metrics
We evaluated the framework based on several key metrics. Narrative coherence was assessed by examining how well the framework maintains the logic of the story. Specifically, we tracked the consistency of character behavior and plot development across episodes to ensure that continuity errors (especially those involving critical content) were detected. Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
III-E Results

### v6

@@ Abstract
Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge. In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies. By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach to identify related episodes and enhance the overall story structure. Experimental results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
Index Terms—  Large language models, AI narrative, story structure.
1 Introduction
Deep learning has transformed multiple domains including NLP, time series analysis and computer vision [1, 2, 3, 4, 5, 6]. Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [7, 8, 9]. However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge [10]. For instance, [11] pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure [12].
@@ 1 Introduction
Fig. 1: Overview of SCORE: (a) Extracts key item statuses per episode. (b) Conducts detailed analysis and summaries of each episode. (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
@@
  • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval. Specially, we outperform baseline GPT model [17] in detecting continuity errors and maintaining overall narrative coherence.
2 Method
@@
2.1 Continuity Analysis and Key Item Status Correction
@@
2.2 Key Item Interaction Analysis
@@
Table 1: Performance Comparison of LLMs with/without SCORE Framework
Model Consistency Coherence Item Status Complex QA
GPT-4 83.21 84.32 0 82.34
SCORE 85.61 ($\uparrow$2.4) 86.9 ($\uparrow$2.58) 98 ($\uparrow$98) 89.45 ($\uparrow$7.11)
GPT-4o 86.78 82.21 0 76.32
SCORE 88.68 ($\uparrow$1.9) 89.91 ($\uparrow$7.7) 96 ($\uparrow$96) 88.75 ($\uparrow$12.43)
Claude3 84.6 80.9 0 69.45
SCORE 87.2 ($\uparrow$2.6) 85.7 ($\uparrow$4.8) 93.1 ($\uparrow$93.1) 83.9 ($\uparrow$14.45)
Gemini-Pro 82.2 83.4 0 78.40
SCORE 85.2 ($\uparrow$3.2) 86.0 ($\uparrow$3.0) 95.0 ($\uparrow$95.0) 85.5 ($\uparrow$7.10)
Qwen-14B 79.3 76.2 0 25.15
SCORE 84.5 ($\uparrow$5.2) 79.1 ($\uparrow$2.9) 86.3 ($\uparrow$86.3) 61.2 ($\uparrow$36.05)
GPT-4o-mini 77.5 75.8 0 23.40
SCORE 81.9 ($\uparrow$4.4) 76.6 ($\uparrow$0.8) 79.8 ($\uparrow$79.8) 60.1 ($\uparrow$36.7)
Llama-13B 71.3 69.8 0 18.72
SCORE 79.1 ($\uparrow$7.8) 73.4 ($\uparrow$3.6) 76.2 ($\uparrow$76.2) 57.3 ($\uparrow$38.58)
2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
@@
For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores [19] between the current episode or query and all other episodes in the vector space. Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$. The top $N$ episodes with the highest $S(e_{c},e_{p})$ scores are retrieved for further analysis, providing a relevant summary of episodes for evaluation or answering questions.
@@
3 Experiments
@@
Table 2: Ablation results for our complete configuration (Full SCORE) versus removing each module.
Configuration Consistency Coherence Item Acc. Complex QA
Full SCORE 89.7 91.2 98.3 88.5
w/o Dynamic Tracking 72.1 ($\downarrow$17.6) 85.4 ($\downarrow$5.8) 61.2 ($\downarrow$37.1) 67.3 ($\downarrow$21.2)
w/o Context Summary 83.2 ($\downarrow$6.5) 68.7 ($\downarrow$22.5) 89.1 ($\downarrow$9.2) 71.4 ($\downarrow$17.1)
w/o Hybrid Retrieval 86.4 ($\downarrow$3.3) 88.9 ($\downarrow$2.3) 94.2 ($\downarrow$4.1) 77.6 ($\downarrow$10.9)
w/o Sentiment 87.1 ($\downarrow$2.6) 89.5 ($\downarrow$1.7) 96.8 ($\downarrow$1.5) 82.3 ($\downarrow$6.2)
Fig. 2: Case study.
3.1 Dataset Preparation
We constructed a story dataset of 5,000 episodes from 1,000 GPT-generated stories evenly distributed across four genres (science fiction, drama, fantasy, comedy). Each story contains 10–15 episodes (avg. 12, $\sim$2,000 tokens), generated using GPT-3.5 and GPT-4 outputs with 15 diverse prompt templates per genre. Quality was ensured via manual filtering by three annotators ($\kappa$=0.78), and complexity stratification included multi-character interactions (30%), non-linear timelines (20%), and symbolic systems (50%). For broader validation, SCORE was also tested on NarrativeQA, BookCorpus, and WP-STORIES, with genre and episode length distributions standardized for cross-dataset comparison. This hybrid setup balances synthetic and human-authored narratives, ensuring robustness across industrial and academic benchmarks.
3.2 Metrics
We evaluated the framework using several key metrics: narrative coherence, assessed by examining logical consistency in character behavior and plot development to avoid continuity errors; consistency, defined as the fraction of responses not conflicting with prior information; coherence, measured as the average semantic or logical flow score across responses; item status, calculated as the proportion of required elements present in each response; and complex question, defined as the fraction of correctly answered complex questions. Finally, we compared the evaluation scores from different methods to measure the stability of the framework.
3.3 Evaluation Process
For each episode, evaluation proceeded in two stages. First, we directly uploaded files to ChatGPT (GPT-4o-mini, GPT-4o, GPT-4) to test baseline narrative assessment without prompts. Second, we performed detailed evaluation using preprocessed files with RAG support, where FAISS-based cosine similarity retrieved semantically and emotionally aligned episodes to provide context. This enriched context was then incorporated into GPT prompts for more accurate key-item tracking and narrative evaluation.
3.4 Baselines
@@
3.5 Main Results